### v1

# Efficient Measure of Information Backflow with Quasi-stochastic Process

Abstract
Characterization and quantification of non-Markovian dynamics in open quantum systems is a topical issue in the rapidly developing field of quantum computation and quantum communication. A standard approach based on the notion of information backflow detects the flow of information from the environment back to the system. Numerous measures of information backflow have been proposed using different definitions of distinguishability between pairs of quantum states. These measures, however, necessitate optimization over the state space which can be analytically challenging or numerically demanding. Here, we propose an alternative witness and measure of information backflow that is explicitly state-independent by utilizing the concept of quasiprobability representation and recent advances in the theory of majorization for quasiprobabilities. We illustrate its use over several paradigmatic examples, demonstrating consistent Markovian conditions with known results and also reported necessary and sufficient condition for qutrit system in random unitary channel. The paper concludes with discussions on the foundational implications of quantum dynamical evolution.
Introduction—In recent years, the study of open quantum systems beyond the standard Markovian approximation has gained considerable interest. In Markovian dynamics, information in a system dissipates continuously to the environment. Interestingly, it has been observed that non-Markovian dynamics exhibit a memory effect such that information can flow back to the system in a revival of the quantum features [1, 2, 3]. This phenomenon has become a subject of interest in the study of decoherence [4, 5], quantum thermodynamics [6], and quantum technologies in general [7]. Thus, an efficient tool to identify and quantify non-Markovianity features is highly desirable.
There have been many proposals for a witness of non-Markovianity with the central object being the quantum dynamical map, $\Lambda_{t}$, $t\geq 0$, which encapsulates the evolution of the open system. In the standard Hilbert space formalism, they are described by completely positive and trace preserving (CPTP) maps that can inform the evolution of a quantum state at any time $t$ via $\rho(t)=\Lambda_{t}[\rho(0)]$. To track information backflow, one is only required to use some notion of distinguishability between two initial states and track its evolution over time. The most natural choice is the trace distance between two states defined as
$D(\rho_{1},\rho_{2})=\frac{1}{2}\left\lVert\rho_{1}-\rho_{2}\right\rVert\,,$ (1)
where $\left\lVert A\right\rVert=\text{Tr}|A|$ is the trace norm of $A$. Due to its contractive property, it is then easy to see that
$D(\Lambda_{t}[\rho_{1}],\Lambda_{t}[\rho_{2}])\leq D(\rho_{1},\rho_{2})\,,$ (2)
where the LHS is a monotonically decreasing function for all time $t\geq 0$ if $\Lambda_{t}$ is Markovian. Therefore, following [8], the BLP (Breuer-Laine-Piilo) measure defined as
[…]
with $\sigma(\rho_{1},\rho_{2};t)\equiv\frac{d}{dt}D(\Lambda_{t}[\rho_{1}],\Lambda_{t}[\rho_{2}])$, quantifies the degree of memory effects. The integral is taken only whenever the information flow $\sigma(\rho_{1},\rho_{2};t)$ is positive — an indication of violating Markovianity. As non-Markovianity is a feature solely due to the dynamics, maximization over all pair states is taken to remove the state dependence.
The above is not the only existing measure of information backflow. Many other proposals utilizing alternative or more general definitions of distances, such as Helstrom matrix [9, 10], and also entropic formulation [11, 12] have been carried out. They all, however, suffer from exact computation as optimizing over state space can be challenging, especially for higher dimensional systems. Even though progress has been made to simplify the optimization task [13], it would be ideal to completely characterize and measure non-Markovianity without this complication. It is also noteworthy to point out that there exists another notion of non-Markovianity based on divisibility [14, 15]. P- or CP-divisibility of a dynamical map concerns whether the intermediate map is still positive or completely positive, respectively, which is then associated with being Markovian. The hierarchical relation between divisibility and information backflow has been studied extensively with the former constituting a stricter criterion for non-Markovianity. Here, we shall only focus on the latter.
In this Letter, we propose a novel measure of non-Markovianity, in particular, information backflow, that requires no optimization component using the framework of quasiprobability representation (QPR) [16, 17]. QPR has been gaining popularity as it provides an alternative mathematical formalism of quantum theories that are reminiscent of probability theory but with the appearance of negative values — widely attributed as the signature of nonclassicality [18, 19, 20, 21, 22, 23]. As such, it has been particularly useful in the study of classical versus quantum theory, identifying resources for quantum advantage, and in general provides new insights about quantum theories. Following the latter point, we will further comment on the implication of our non-Markovianity witness in a quantum foundational perspective. We note that previous work [24] had considered the use of frame representation to develop a new measure. However, similar to the previously mentioned measures, it also involves maximizing over all possible pairs of initial states. Crucially, their work focuses on studying non-Markovianity in continuous variable system, but here we focus on the discrete case.
Quasiprobability representation—In QPR, elements of quantum theories, such as states and unitaries, are described by vectors or matrices consisting of real values that typically obey a normalization constraint. To translate quantum objects defined in the standard Hilbert space formalism into its QPR, the mapping is achieved by the so-called frame and its dual frame [16, 17]. Hence, it is also colloquially known as frame representation. In the following, we will introduce the mathematical formalism of QPR.
For a system in Hilbert space $\mathcal{H}^{d}$, a frame $\{F_{j}\}$ is defined as a set of operators that span the Hermitian space. That is, it forms, in general, an overcomplete basis for the space of all Hermitian operators acting on $\mathcal{H}^{d}$. For a given frame, the dual frame $\{G_{j}\}$ is also a complete basis of a Hermitian space that satisfy $A=\sum_{j}\text{Tr}\left[AF_{j}\right]G_{j}$ for all Hermitian operator $A$. Note that the dual frame is generally not unique except for the case when the frame is minimal, i.e., |$|\{F_{j}\}|=d^{2}$. In this case, they obey the relation $\text{Tr}\left[F_{j}G_{k}\right]=\delta_{jk}$. The QPR of a state $\rho$, channel $\mathcal{E}$, and measurement effect $M$ are then obtained via
$\rho\longrightarrow q^{\rho}_{j}$ $=$ $\text{Tr}\left[\rho F_{j}\right]\,,$ (4)
$\mathcal{E}\longrightarrow S^{\mathcal{E}}_{jk}$ $=$ $\text{Tr}\left[F_{j}\mathcal{E}[G_{k}]\right]\,,$ (5)
$M\longrightarrow v^{M}_{j}$ $=$ $\text{Tr}\left[MG_{j}\right]\,.$ (6)
Note that the above always return real values and they correspond to the entries of the column vector $\mathbf{q}^{\rho}=[q^{\rho}_{j}]$, square matrix $S^{\mathcal{E}}=[S^{\mathcal{E}}_{jk}]$, and row vector $\mathbf{v}^{M}=[v^{M}_{j}]$. From here, the Born rule probability of measuring a state undergoing an evolution beforehand is simply calculated following vector/matrix operation:
$P(M|\rho,\mathcal{E})=\text{Tr}\left[M\mathcal{E}[\rho]\right]$ $=$ $\mathbf{v}^{M}S^{\mathcal{E}}\mathbf{q}^{\rho}$ (7)
$=$ $\sum_{jk}v^{M}_{j}S^{\mathcal{E}}_{jk}q^{\rho}_{k}\,.$
Here, we consider a normalization condition on the frame operators, $\sum_{j}F_{j}=\mathbb{1}$. This implies that $\sum_{j}q^{\rho}_{j}=1$ so it can interpreted as a quasiprobability distribution. Moreover, we also imposed that $\text{Tr}\left[G_{j}\right]=1\,\forall\,j$ to ensure that summing up the effects in a POVM (positive operator-value measure) returns a unit measurement: $\sum_{k}v^{M_{k}}_{j}=1\,\forall j$, where $\{M_{k}|M_{k}\geq 0,\sum_{k}M_{k}=\mathbb{1}\}$ is a POVM. As a result, $v^{M_{k}}_{j}$ can be thought of as a conditional quasiprobability distribution. Lastly, based on the above condition, one also finds that $\sum_{j}S^{\mathcal{E}}_{jk}=1\,\forall k$, which means that it is now a quasi-stochastic matrix. If the channel is unital, i.e., $\mathcal{E}[\mathbb{1}]=\mathbb{1}$, then $S^{\mathcal{E}}$ is quasi-bistochastic, [^1] In some literature, ‘bistochastic’ is also referred as ‘doubly stochastic’. $\sum_{j}S^{\mathcal{E}}_{jk}=\sum_{k}S^{\mathcal{E}}_{jk}=1$. As can be seen, the elements of quantum theories have been described in a language that is similar to classical theories and hence is natural to use for the study of classical-quantum distinction.
Majorization theory for quasiprobability distributions—To derive the main results of this paper, we shall make use of the observations made in [25]. There, the authors extended the theory of majorization for quasiprobabilities and shown that certain family of entropries are well-defined and meaningful even for quasiprobabilities. Of particular interest, Theorem 9 of [25] states that:
If $\alpha=\frac{2a}{2b-1}$ for positive integers $a,b$ with $a\geq b$, then $H_{\alpha}(\mathbf{q})$ is well-defined on the set of quasi-distributions, and if $\mathbf{q}\succ\mathbf{q}^{\prime}$ for two quasi-distributions $\mathbf{q},\mathbf{q}^{\prime}$ then $H_{\alpha}(\mathbf{q})\leq H_{\alpha}(\mathbf{q}^{\prime})$.
Here, $H_{\alpha}(\mathbf{q})$ is the Renyi-$\alpha$ entropy [26] defined as
$H_{\alpha}(\mathbf{q})=\frac{1}{1-\alpha}\log\left[\sum_{k}q_{k}^{\alpha}\right]\,,$ (8)
and $\mathbf{q}\succ\mathbf{q}^{\prime}$ denotes that $\mathbf{q}$ majorizes $\mathbf{q}^{\prime}$ in the preorder sense such that there exists a bistochastic map $A$ that satisfies the relation $A\mathbf{q}=\mathbf{q}^{\prime}$.
The result above establishes the Schur-concavity property and monotonicity of Renyi-$\alpha$ entropy on the set of quasiprobability distributions under a subset of $\alpha$ values. Moreover, the choice of $\alpha$ also ensures that the function is real-valued and satisfies many of the properties that Renyi originally invoked [27].
Main results—In the following, we shall take advantage of the result above to derive a new witness and measure of non-Markovianity. Let us make the choice of $\alpha=2$ for the Renyi-$\alpha$ entropy. This form is also known as collision entropy, and its usage has been explored previously as entropy of Wigner function [28], in relation to uncertainty principle [29], and in axiomatization of quantum theories [30, 31, 32].
Considering state transforming over time given by $\mathbf{q}(t)=S^{\Lambda_{t}}\mathbf{q}(0)$, where we denote $\mathbf{q}\equiv\mathbf{q}^{\rho}$ for brevity, then it follows from the above that
$H_{2}(\mathbf{q}(t))\geq H_{2}(\mathbf{q}(0))$ (9)
if $S^{\Lambda_{t}}$ is bistochastic for all time $t\geq 0$. One can then immediately show that under Markovian condition the monotonicity holds, and violation occurs for non-Markovian case. This formulation, however, is akin to the entropic-based formulation in [11] and would still entail a maximization over the state space that is inefficient.
Surprisingly, we show that it is possible to further simplify the above relation and remove the dependence on state. Notice that the Renyi-2 entropy can be expressed more compactly as
$H_{2}(\mathbf{q})=-\log\left[\sum_{k}q_{k}^{2}\right]=-\log\mathbf{q}^{T}\mathbf{q}\,,$ (10)
where $\mathbf{q}^{T}$ is the transpose of $\mathbf{q}$ into a row vector. Therefore, Eq. (9) can be written as $-\log[(S^{\Lambda_{t}}\mathbf{q})^{T}(S^{\Lambda_{t}}\mathbf{q})]=-\log[\mathbf{q}^{T}(S^{\Lambda_{t}})^{T}S^{\Lambda_{t}}\mathbf{q}]\geq-\log\mathbf{q}^{T}\mathbf{q}$ and can be simplified neatly as
$(S^{\Lambda_{t}})^{T}S^{\Lambda_{t}}\leq\mathbb{1}\,.$ (11)
Under Markovian condition, the quasi-stochastic map of a channel $\Lambda_{t}$ must then obey this monotonic relation. More explicitly, the eigenvalues of $(S^{\Lambda_{t}})^{T}S^{\Lambda_{t}}$ must be a monotonically decreasing function over time, and the violation of the monotonicity signifies non-Markovianity. As such, in the same spirit as previously proposed measures, we define our measure of non-Markovianity as
[…]
$\text{with}\quad\zeta(t)$ $=$ $\frac{d}{dt}\left\lVert(S^{\Lambda_{t}})^{T}S^{\Lambda_{t}}\right\rVert\,.$ (12)
We remark that this formulation is reminiscent of the measure for P- or CP-divisibility [14] such that it is a sole function on the dynamical map so no optimization over state is required and the eigenvalues of the map obey a monotonic relation.
Let us now discuss some of the properties of the above measure. Firstly, we have that $(S^{\Lambda_{t}})^{T}=S^{\Lambda_{t}^{\dagger}}$. From here, it is easy to see that the equality in Eq. (11) is saturated if the channel is unitary, which reduces into orthogonality relation. This is obvious from the basic property of unitary matrices: $U^{\dagger}=U^{-1}$. Interestingly, it has been shown before that an axiom based on the orthogonality of the quasi-bistochastic map enable for the reconstruction of qubit and its dynamics [32].
Secondly, an equivalent witness can be obtained by swapping the matrices in Eq. (11), i.e, $S^{\Lambda_{t}}(S^{\Lambda_{t}})^{T}\leq\mathbb{1}$. This is clear as the eigenvalues of $AB$ is equal to $BA$. Moreover, both instances of the product matrix yield a symmetric matrix. The observation on the symmetry of the witness has a foundational implications that will be discussed later.
Next, we shall see our measure of non-Markovianity in action over several paradigmatic examples with qubit and qutrit systems. Notice that the measure above is derived without any reference for a specific frame representation. Thus, the result works in general for any choice of quasiprobability representation. In order to make specific calculation, here we pick the discrete Wigner representation with Wootters’ representation for the qubit case [33] and Gross’ for the qudit case (odd prime dimension) [34]. In this representation, they belong to the family of normal quasiprobability representation as their frame and dual frame operators are proportional to each other [35]. In fact, they are related via $G_{j}=dF_{j}$ since the frame is minimal. The Wootters’ discrete Wigner representation has frame operators given by
$F_{j}=\frac{1}{4}\left[\mathbb{1}+(-1)^{j_{1}}\sigma_{z}+(-1)^{j_{2}}\sigma_{x}+(-1)^{j_{1}+j_{2}}\sigma_{y}\right]\,,$ (13)
where $j\equiv(j_{1},j_{2})\in\mathbb{Z}_{2}\times\mathbb{Z}_{2}$, and $\sigma_{x},\sigma_{y},\sigma_{z}$ are qubit Pauli operators. The construction for discrete Wigner representation is more involved and since we are only working with qutrit system in the example below we show the frame operators for $d=3$ in the Appendix.
We ascertain that the results below are frame independent by repeating the calculation with other representation, namely, the symmetric informationally-complete (SIC)-POVM representation [36].
Example 1: Pure decoherence model—Let us start with the prototypical model of a two-level system in a bosonic bath environment undergoing a pure decoherence dynamics [11]. The evolution of the system’s state is given by the master equation
$\mathcal{L}[\rho(t)]=\frac{d}{dt}\rho(t)=\gamma(t)\left[\sigma_{z}\rho(t)\sigma_{z}-\rho(t)\right]\,,$ (14)
where the time-dependent decay rate $\gamma(t)=-\frac{1}{G(t)}\frac{d}{dt}G(t)$ with $G(t)$ being the decoherence function. From this, the dynamical map reads
$\Lambda_{t}:\rho(0)=\begin{bmatrix}\rho_{00}&\rho_{01}\\ \rho_{10}&\rho_{11}\end{bmatrix}\longrightarrow\rho(t)=\begin{bmatrix}\rho_{00}&G(t)\rho_{01}\\ G(t)\rho_{10}&\rho_{11}\end{bmatrix}\,.$ (15)
Note that the map $\Lambda_{t}$ is CPTP for $G(t)\leq 1$.
To find the non-Markovianity criteria, we first find the quasi-stochastic representation of $\Lambda_{t}$ following the Wootters’ representation:
$S^{\Lambda_{t}}=\frac{1}{2}\begin{bmatrix}1+G&1-G&0&0\\ 1-G&1+G&0&0\\ 0&0&1+G&1-G\\ 0&0&1-G&1+G\end{bmatrix}\,,$ (16)
where $G\equiv G(t)$ for simplicity. Notice that here $S^{\Lambda_{t}}$ is self-transpose, hence $\Lambda_{t}$ is its self-dual, i.e., $\Lambda_{t}^{\dagger}=\Lambda_{t}$. More importantly, the quasi-bistochastic map is represented non-negatively (same with SIC-POVM representation) that affords us to use the main result in Eq. 9. In fact, all the examples in this paper have a map that are non-negative. Calculating the LHS of Eq. (11), we find its eigenvalues are given by $\{1,1,G^{2},G^{2}\}$. Using Eq. (Efficient Measure of Information Backflow with Quasi-stochastic Process), we obtain that
$\zeta(t)=4G(t)\frac{d}{dt}G(t)=-4\gamma(t)G^{2}(t)\,.$ (17)
[…]
Example 2: Dissipation model—Next, we consider the same system and bath environment but with system-environment interaction that allows for dissipation in the populations too. This evolution is governed by the time-local master equation
$\frac{d}{dt}\rho(t)=$ $-\frac{i}{2}s(t)[\sigma_{+}\sigma_{-},\rho(t)]$ (18)
$\quad+\gamma(t)\left[\sigma_{-}\rho(t)\sigma_{+}-\frac{1}{2}\big{\{}\sigma_{+}\sigma_{-},\rho(t)\big{\}}\right]\,,$
where $\sigma_{\pm}=(\sigma_{x}\pm i\sigma_{y})/2$ is the raising and lowering operators, $s(t)=-2\text{Im}[\dot{G}(t)/G(t)]$ is the time-dependent Lamb shift, and $\gamma(t)=-2\text{Re}[\dot{G}(t)/G(t)]=-\frac{2}{|G(t)|}\frac{d}{dt}|G(t)|$ is the decay rate. The dynamical map can be shown to be
$\Lambda_{t}:\rho(0)\rightarrow\rho(t)=\begin{bmatrix}\rho_{00}+(1-|G(t)|^{2})\rho_{11}&G(t)^{*}\rho_{01}\\ G(t)\rho_{10}&|G(t)|^{2}\rho_{11}\end{bmatrix}\,.$ (19)
Following the same calculation steps as the previous example, we find the eigenvalues of $(S^{\Lambda_{t}})^{T}S^{\Lambda_{t}}$ as $\{|G|^{2},|G|^{2},1-|G|^{2}+|G|^{4}-\sqrt{(-1+|G|^{2})^{2}(1+|G|^{4})},1-|G|^{2}+|G|^{4}+\sqrt{(-1+|G|^{2})^{2}(1+|G|^{4})}\}$. From here, we find
$\zeta(t)=8|G(t)|^{3}\frac{d}{dt}|G(t)|=-4\gamma(t)|G(t)|^{4}\,,$ (20)
and identify that non-Markovianity is observed when $\gamma(t)<0$. Again, the witness here coincides with the BLP and CP-divisibility measure [1].
Example 3: Random unitary channel—In the last example, we shall look into evolution under random unitary channel of a qubit and qutrit system. The general qudit system has been studied in [37]. Here, we consider the master equation
$\frac{d}{dt}\rho(t)=\sum_{\alpha=1}^{d^{2}-1}\gamma_{\alpha}(t)\left[U_{\alpha}\rho(t)U_{\alpha}^{\dagger}-\rho(t)\right]\,,$ (21)
where $U_{\alpha}\equiv U_{k,l}=\sum_{m=0}^{d-1}\omega^{mk}|m\rangle\langle m+k|$ is the Weyl (generalized spin) operator with $\alpha\equiv(k,l)$ via $\alpha=kd+l$, $k,l\in\mathbb{Z}_{d}$, and $\omega=e^{i2\pi/d}$. The dynamical map takes the form
$\Lambda_{t}[\rho]=\sum_{\alpha=0}^{d^{2}-1}p_{\alpha}(t)U_{\alpha}\rho U_{\alpha}^{\dagger}\,,$ (22)
where
$p_{\alpha}(t)$ $=$ $\frac{1}{d^{2}}\sum_{\beta=0}^{d^{2}-1}H_{\alpha\beta}\lambda_{\beta}(t)\,,$
$\lambda_{\beta}(t)$ $=$ $\exp\left(\sum_{k=1}^{d^{2}-1}H_{\beta k}\Gamma_{k}(t)\right)\,,$ (23)
[…]
For $d=2$, the Weyl operators reduce to the identity and Pauli operators. In this case, we find the criterion for Markovian condition to be
$\gamma_{1}(t)+\gamma_{2}(t)$$\geq$ $0\,,$
$\gamma_{1}(t)+\gamma_{3}(t)$$\geq$ $0\,,$
$\gamma_{2}(t)+\gamma_{3}(t)$$\geq$ $0\,.$ (24)
This coincides with the BLP and P-divisibility condition as expected [38] but differs from the CP-divisibility condition, which is $\gamma_{k}(t)\geq 0$. This is an instance of the departure of the Markovian condition between information backflow and divisibility with the former being weaker.
In $d=3$, we find the Markovian criterion as
$\gamma_{1}(t)+\gamma_{2}(t)+\gamma_{4}(t)+\gamma_{5}(t)+\gamma_{7}(t)+\gamma_{8}(t)$$\geq$ $0\,,$
$\gamma_{1}(t)+\gamma_{2}(t)+\gamma_{3}(t)+\gamma_{4}(t)+\gamma_{6}(t)+\gamma_{8}(t)$$\geq$ $0\,,$
$\gamma_{1}(t)+\gamma_{2}(t)+\gamma_{3}(t)+\gamma_{5}(t)+\gamma_{6}(t)+\gamma_{7}(t)$$\geq$ $0\,,$
$\gamma_{3}(t)+\gamma_{4}(t)+\gamma_{5}(t)+\gamma_{6}(t)+\gamma_{7}(t)+\gamma_{8}(t)$$\geq$ $0\,.$ (25)
To our knowledge, characterization based on BLP and P-divisibility appears to be highly non-trivial and only sufficient (but not necessary) condition had been obtained:
$\gamma_{i_{1}}(t)+\gamma_{i_{2}}(t)+\gamma_{i_{3}}(t)\geq 0$ (26)
for all triples $\{i_{1},i_{2},i_{3}\}\subset\{1,\dots,8\}$. Similar to the qubit case, the evolution is CP-divisible if all rates are non-negative, $\gamma_{\alpha}(t)\geq 0$.
Discussions—The monotonic relation in Eq. (11) provides an interesting perspective on non-Markovianity between classical and quantum theory. In the classical case, evolution is generally described by a stochastic map. Bistochastic map, a subset of all stochastic maps, has a special property such that its transpose returns a valid map as well. One can think of the transpose as an attempt to ‘reverse’ the transition probabilities between the input and output. Recent works [39, 40, 41] have also pointed out that Hermitian adjoint and transpose of a quantum and classical process, respectively, correspond to the time-reversal of the process. Interestingly, under QPR the time-reversal notion has been unified for both theories with the transpose operation. Reversal of the process, however, does not always help in reversing the state of the system. This is also reflected in the monotonic relation, in both forms of $S^{T}S$ and $SS^{T}$, that it will mix any probability distribution further. In contrast, applying forward and reverse quantum processes successively, or vice versa, on any state may result in a momentary revival of quantum features albeit still leading to noise eventually.
It should be noted that ‘reversal’ is not equivalent to the inverse, which is generally not a valid map except for unitaries. [^2] For classical theory, unitary processes are described by permutation matrices — matrices with a single $1$ in each row and columns and 00 elsewhere Nevertheless, for an irreversible process one can still construct a map that approximately invert the forward process known as Bayesian inference. Remarkably, the canonical quantum Bayesian inference, Petz map, also contains $S^{T}$ which might imply that it is a useful ingredient in retrieving loss information [42].
In formulating the condition for Markovianity, it is also required that quantum maps are represented non-negatively. This is seemingly counterintuitive to current perception where it is widely believed that the appearance of negativity is necessary for a lot of quantum phenomena. Thus, one would also think that quantum non-Markovianity, a non-classical phenomenon, would occur due to negativity in the dynamical map. It would be interesting to show if a negative quasi-stochastic map corresponding to a Markovian evolution would still satisfy the monotonic relation in Eq. (9). Alternatively, it would also be interesting if it is always possible to find a frame representation that will return a non-negative quasi-stochastic matrix.
Although the QPR of the dynamical maps studied above are represented non-negatively, we would like to point out that the generator $\mathcal{L}$ has a non-classical representation. More explicitly, we find $L=[L_{jk}]$, $L_{jk}=\text{Tr}\left[F_{j}\mathcal{L}[G_{k}]\right]\in\mathbb{R}$, which is a $d^{2}\times d^{2}$ matrix that satisfies $\sum_{j}L_{jk}=0\,\forall\,k$. This is the quantum analogue of the classical Kolmogorov generator [43] that additionally satisfies $L_{jk}\geq 0$ for $j\neq k$, i.e., its off-diagonal elements are non-negative. For the pure decoherence and random unitary channel models, we observe that the off-diagonal entries of $L$ become negative when non-Markovianity condition is satisfied and non-negative otherwise. However, for dissipative model the matrix $L$ has negative non-diagonal entries in general. From here, it suggests that negativity in the generator appears to be necessary but not sufficient for non-Markovianity. This observation has also been pointed out in [44], where they studied its connection to classical and quantum P-divisibility. It remains unclear how exactly negativity in the generator of quantum evolution contributes to the degree of non-Markovianity. We believe that the study using QPR would be crucial in gaining a fundamental understanding of the difference between classical and quantum processes.
Conclusions – In this Letter, we have proposed a new witness and measure of non-Markovianity using the formalism of quasiprobability representation. In particular, the measure detects information backflow through the monotonic relation on the product of quasi-bistochastic map, representing the dynamical map, and its transpose. The measure is explicitly dependent only on the dynamical map, thus foregoing any optimization task that plagues much of the previously proposed measure. This is particularly useful when studying evolution of a quantum system of large dimensions. We illustrate the use over several paradigmatic examples with explicit calculation done in the discrete Wigner and SIC-POVM representation — the two most widely used frame representations for discrete quantum mechanics. We further comment on the non-classicality of the evolution and believe that the process’ generator, being the one possessing the negativity, should be a focus for future work in its role in quantum non-Markovianity.
Acknowledgements.—We thank Dagomir Kaszlikowski and Lim Zi Yao for discussions. This project was supported by the Ministry of Education, Singapore, under its Academic Research Fund Programme (RG154/24). K.O. was supported by grants from the National Research Foundation, Singapore and A*STAR under its Quantum Engineering Programme 2.0 (NRF2021-QEP2-02-P07) and the Singapore Ministry of Education Academic Research Fund (T2EP50222-0038).
Appendix: Discrete Wigner representation for qutrit—The following are the frame operators for the discrete Wigner representation in $d=3$:
$F_{00}=\frac{1}{3}\begin{bmatrix}1&0&0\\ 0&0&1\\ 0&1&0\end{bmatrix}\,,\quad$ $F_{01}=\frac{1}{3}\begin{bmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{bmatrix}\,,$ (27)
$F_{10}=\frac{1}{3}\begin{bmatrix}0&1&0\\ 1&0&0\\ 0&0&1\end{bmatrix}\,,\quad$ $F_{10}=\frac{1}{3}\begin{bmatrix}1&0&0\\ 0&0&-\omega\\ 0&\omega^{2}&0\end{bmatrix}\,,$
$F_{11}=\frac{1}{3}\begin{bmatrix}0&0&\omega^{2}\\ 0&1&0\\ -\omega&0&0\end{bmatrix}\,,\quad$ $F_{12}=\frac{1}{3}\begin{bmatrix}0&-\omega&0\\ \omega^{2}&0&0\\ 0&0&1\end{bmatrix}\,,$
$F_{20}=\frac{1}{3}\begin{bmatrix}1&0&0\\ 0&0&\omega^{2}\\ 0&-\omega&0\end{bmatrix}\,,\quad$ $F_{21}=\frac{1}{3}\begin{bmatrix}0&0&-\omega\\ 0&1&0\\ \omega^{2}&0&0\end{bmatrix}\,,$
$F_{22}=\frac{1}{3}\begin{bmatrix}0&\omega^{2}&0\\ -\omega&0&0\\ 0&0&1\end{bmatrix}\,,\quad$
where $\omega=e^{i2\pi/3}$ is the cube root of unity.